# Perseus: Removing Energy Bloat from Large Model Training

### v1

@@ 6.4 Iteration Time–Energy Frontier Comparison
Figure 9: Iteration time–energy frontiers for GPT-3, achieved by Perseus and the two baselines derived from Zeus [68]. Perseus Pareto-dominates all other approaches. The dotted vertical line is the iteration time of running all GPUs at their maximum power limit, which is the default mode of operation. Please note the different Y-axes.
@@
The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm. Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus [68], which is an energy optimization framework for a single-GPU training with the training time–energy frontier. We implemented two Zeus-based baselines to make it work in parallelization configurations and generated “iteration time–energy” frontiers.
  1. 1.
@@
  2. 2.
    ZeusPerStage: Sets one power limit per stage that balances forward computation time.
We run training and measurement under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) hybrid parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40. Figure 9 shows the frontiers of Perseus and Zeus for different sizes of GPT-3 under three parallelization configurations. Frontiers for other models are in Appendix H.
Perseus Pareto-dominates Zeus. ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage equally, unable to reduce intrinsic energy bloat. While ZeusPerStage can balance the forward computation time of each stage, it is unaware of the critical path of the DAG, slowing down critical computations. In contrast, Perseus can precisely slow down non-critical computations, tightly packing computation.
6.5 Overhead of Perseus
@@
Our online profiling method (§5) introduces extra training time by running some iterations at a lower frequency. For our A100 workloads, the average percentage of slowdown for such iterations was 8.2%, and the average extra training time introduced by profiling was 13 minutes. This is negligible compared to how long large model training can take.
@@
The average time spent on running Perseus’s optimization algorithm (§4) across the five A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes). For our largest scale emulation experiment (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds. While the runtime of the algorithm will increase with larger DAGs for larger models, we believe the overhead be justified because training time also increases with the scale of the training job. Finding the energy-optimal iteration time and energy schedule given the straggler’s iteration time for extrinsic energy bloat reduction is instant.
7 Related Works
@@
Many recent works focus on enabling and accelerating large model training using 3D parallelism (data, tensor, and pipeline parallelism). GPipe [26] and PipeDream [41] were the first to introduce pipeline parallelism and explicitly discussed the difficulty of perfectly balancing computation time across stages. Megatron-LM [56, 43] is a Transformer-based large model training framework that provides efficient manually designed execution plans. Later on, modern training frameworks that focus on large scale training [53, 59, 36, 37, 40, 34] were introduced to support various models at scale. DeepSpeed [53] introduced ZeRO redundancy optimizer that shards model states, which is recently widely used for large model training [52, 72]. Alpa [73] and GSPMD[70] are automatic parallelization framework for general DNNs. Finally, some recent works have looked into fault-tolerant large model training as well [61, 27, 9]. Unfortunately, energy consumption is not an optimization metric for any of the major large model training frameworks.
@@
A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training, including those that present observations and estimations [48, 60, 33, 16, 38] and those that propose optimization methods for training time, energy consumption, and carbon footprint [74, 63, 68, 67, 32, 13].
Zeus [68] is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on simple single-GPU training. EnvPipe [13], on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown. However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings and an infinite loop during optimization. Perseus Pareto-dominates both Zeus and EnvPipe by viewing large model training as a computation DAG and introducing a principled optimization algorithm. Perseus is also the first to introduce the notion of extrinsic energy bloat in large model training and optimize both simultaneously.
8 Conclusion
We presented Perseus, an energy optimization system for large model training. Perseus builds on top of observation that there are fundamental computation imbalance at different levels in large model training that causes intrinsic and extrinsic energy bloat. We introduced a principled graph cut-based algorithm that simultaneously reduces both.
Perseus advances the state-of-the-art of DNN training energy optimization by establishing a new time–energy Pareto frontier for large model training. The integration of Perseus into the training workflow has strong implications for the future of AI development. Importantly, reducing energy bloat leads to practically no latency and throughput degradation, which has the potential to greatly enhance the sustainability of distributed training in the midst of recent proliferation of LLMs and GenAI.
Acknowledgements
We thank Chameleon Cloud [30] for providing A100 nodes.
Appendix A Workload Details
A.1 Minimum Imbalance Pipeline Partitioning
@@
[…]
Table 6: Imbalance ratio between the longest and the shortest stages for various models. 1.00 would mean perfect balance. Partitions for $N$ stages is a list of $N+1$ numbers, where the numbers represent layer indices. For instance, [0, 6, 12, 19, 25] for GPT-3 1.3B means there are 6, 6, 7, and 5 Transformer layers in each stage, and the final stage also has the language model head.
@@
Arguably, these are one of the most homogeneous large models because they are a stack of identical Transformer [62] encoder or decoder layers. However, the very last layer is the language modeling head, which maps the final features to probabilities over the entire vocabulary. The vocabulary size of GPT-3 is 50k, Bloom 251k, and BERT 31k, which results in a very large linear layer for the last stage. This leads to varying amounts of imbalance and different minimum imbalance partitioning results for each models.
@@
A.2 Experiment Parameters
@@
Table 7: Experiment Parameters for 3D parallelism experiments on A40 GPUs. Microbatch size is per-pipeline, and there are two data parallel copies of the same pipeline. Thus, global batch size should be calculated as the product of microbatch size and the number of microbatches times two.
@@
Table 8: Experiment Parameters for eight-stage pipeline parallelism experiments on A40 GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
@@
Table 9: Experiment Parameters for Pipeline Parallelism Experiments on A100 PCIe GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
Tables 78, and 9 list model variant names and experiment parameters for our experiments. Model names and configurations for GPT-3 were taken as is from the original model publication [11]. Especially, model names and configurations for BERT and T5 were directly taken from the Huggingface Hub pretrained model zoo, except for the huge variant of BERT, which we created to have hidden dimension 2048. Wide-ResNet was based on Torch Vision [45] but scaled up following the model’s original publication [69] using its width factor parameter. The unit time parameter $\tau$ was set to 1 ms for all experiments.
Appendix B Pipeline Energy Minimization is NP-Hard
The original Pipeline Energy Minimization problem is stated in formal terms:
$\min_{F}\quad$ $\textrm{Energy}(F)$ (6)
@@ Appendix B Pipeline Energy Minimization is NP-Hard
B.1 One Stage Two Frequencies Simplification
@@
B.3 NP-Hardness Proof
Theorem 2.
@@ B.3 NP-Hardness Proof
Appendix C Visualizations for Intrinsic Energy Bloat
Figure 10: Visualization of Perseus’s solution for four stage pipeline workloads on NVIDIA A100 PCIe GPUs. For each workload, the left is running every computation at maximum frequency, and the right is Perseus’s energy schedule that reduces only intrinsic energy bloat without inflating iteration time. Note that these are not real workloads we run in Section 6; real workloads have far more microbatches.
Appendix D Full Details of ReduceTime
@@ D.1 Generating Capacity DAG
$\min_{I^{+},\leavevmode\nobreak\ I^{-}}\quad\sum_{i\in I^{+}}{e_{i}^{+}}-\sum_{i\in I^{-}}{e_{i}^{-}},$ (7)
where $e_{i}(t_{i})$ is the exponential function fit to Pareto-optimal (time, energy) measurments for computation $i$, $e_{i}^{+}=e_{i}(t_{i}-\tau)-e_{i}(t_{i})$ is the extra amount of energy needed to speed up $i$ by $\tau$ and $e_{i}^{-}=e_{i}(t_{i})-e_{i}(t_{i}+\tau)$ is the energy saved from slowing down $i$ by $\tau$.
@@
$(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$ (8)
@@
D.2 Max Flow Algorithm on the Capacity DAG
@@
3
4
Construct a new graph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding new source and sink nodes $s^{\prime}$ and $t^{\prime}$, edges from $s^{\prime}$ to each node in $V$, edges from each node in $V$ to $t^{\prime}$, and an edge from $t$ to $s$triangleright Define the capacity $c^{\prime}(e)$ of each edge $e\in E^{\prime}$
5: for $v\in V$ do
6       $c^{\prime}(s^{\prime}v)\leftarrow\sum_{u\in V}l(uv)$ $c^{\prime}(vt^{\prime})\leftarrow\sum_{w\in V}l(vw)$
@@ 9       $c^{\prime}(uv)\leftarrow u(uv)-l(uv)$
$c^{\prime}(ts)\leftarrow\infty$triangleright Find the max flow on G’
$f^{\prime}\leftarrow$ EdmondsKarp ($G^{\prime}$, $c^{\prime}$) $\triangleright$ $G$ has a feasible flow if and only if $G^{\prime}$ has a saturating flow
@@ Appendix E Proof for Polynomial Runtime
$t_{\max}$ and $t_{\min}$ can be constructed by multiplying the number of computations with the average execution time of an computation. Computations are executed with frequencies $f_{\min}$ and $f_{\max}$, respectively, and thus the multipliers of $N$ and $M$ do not cancel out when $t_{\max}-t_{\min}$ is evaluated. Therefore, $t_{\max}-t_{\min}$, and hence $(t_{\max}-t_{\min})/\tau$, is $O(N+M)$. ∎
Appendix F Underestimations of the Emulator
Model Leftmost point Rightmost point
gpt3-xl 20.5% 23.6%
bert-huge-uncased 26.7% 35.3%
t5-3b 19.1% 19.1%
bloom-3b 9.3% 10.7%
wide-resnet101 (width factor 8) 17.5% 19.8%
Table 10: By how much the emulator underestimates the percentage of energy savings for four-stage pipeline workloads on A100 PCIe. Leftmost point represents intrinsic energy bloat, whereas the rightmost point extrinsic.
Our emulator, used to present energy saving results for large scale workloads (GPT-3 175B and Bloom 176B), always underestimates energy savings compared to actual measurements, and argue that the emulated results are empirically a lower bound of actual savings. The characteristics of time and energy changes depends more on the type of the GPU than the parallelization setting as different GPU generations show drastically different sensitivity in terms of time and energy when frequency is varied. Therefore, we compare results from the A100 PCIe GPU.
Table 10 shows the percentage of energy savings underestimation for two points on the “iteration time–energy” frontier:
  • Leftmost point: The point with the shortest iteration time, corresponding to how much the emulator underestimates the amount of possible intrinsic energy bloat reduction.
  • Rightmost point: The point with the longest iteration time, corresponding to how much the emulator underestimates the amount of possible energy savings brought about by slowing down the entire pipeline in the case of a straggler.
We believe that underestimation is happening due to our simplifying assumption that $P_{\mathrm{blocking}}$, the power consumption of the GPU while blocking on communication, is constant regardless of the GPU’s frequency. In reality, $P_{\mathrm{blocking}}$ would be an increasing function of frequency. Therefore, when the emulator assumes a constant $P_{\mathrm{blocking}}$ when calculating the energy consumption of a pipeline that is overall running on a reduced frequency, it will be overestimating its total energy consumption, leading to underestimations.
Appendix G Large-Scale Emulation
Section 6.3 used A100 SXM GPUs, which we momentarily had access to, in order to profile model layers and run large scale emulation. This is because we believe that SXM models are more representative of real large scale GPU clusters that train large models. Nonetheless, in this section, we provide the same set of results from A100 PCIe GPUs, as well, along with the same A40 GPUs results. All other experiment methodology and parameters are identical.
Results Summary.
Figure 11: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1024 GPUs.
Figure 11 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1024 GPUs as a representative.
Intrinsic Bloat Reduction.
Model GPU Type
Energy Savings (%)
given # Microbatches
12 24 48 96
GPT-3 175B A100 9.02 9.77 10.17 10.39
A40 11.81 10.22 9.34 8.88
Bloom 176B A100 6.39 4.99 4.21 3.80
A40 6.97 4.49 3.12 2.41
Table 11: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
Table 11 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches. GPT-3 175B on A100 is unique and has the opposite trend compared to others: intrinsic energy bloat reduction increases with the number of microbatches. The opposite trends stem from the amount of stage imbalance on each GPU, and the amount of slowdown the minimum-energy frequency incurs. What is unique about GPT-3 175B on A100 is that the amount of imbalance is large whereas the minimum-energy frequency incurs little slowdown for forward and backward computations. Therefore, only for GPT-3 175B on A100, microbatches in the middle of the pipeline are able to fully slow down to their potential. Thus, when such microbatches are added to the pipeline, the percentage of intrinsic energy bloat reduction increases.
Extrinsic Energy Bloat.
Figure 12: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines. When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$. More blocking time after $T^{*}$ decreases overall energy savings. Please note the different Y-axes.
Figure 12 presents extrinsic energy bloat results in the strong scaling configuration, where a straggler is slower than others in varying degrees of slowdown from 1.05 to 1.50.
Appendix H Iteration Time–Energy Frontiers

### v2

@@ 6.4. Iteration Time–Energy Frontier Comparison
Figure 9. [Experiment] Iteration time–energy frontiers for GPT-3, achieved by Perseus and the two baselines derived from Zeus (zeus-nsdi23, ). Perseus Pareto-dominates all other approaches. The dotted vertical line is the iteration time of running all GPUs at their maximum power limit, which is the default mode of operation. Please note the different X- and Y-axes.
@@
The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm (§4). Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus (zeus-nsdi23, ). Since Zeus only produces the training time–energy frontier for single-GPU training jobs, we implemented two Zeus-based baselines for large model training scenarios:
@@
  • ZeusPerStage: Finds a set of per-stage power limits that balances forward computation time.
Figure 9 shows the frontiers of all solutions for different sizes of GPT-3 under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) 3D parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40. All results were obtained from running on testbed GPUs. Appendix G has frontiers for other models.
Perseus Pareto-dominates Zeus. ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, unable to reduce intrinsic energy bloat. ZeusPerStage can balance the forward computation time of each stage, but is unaware of the critical path of the DAG, slowing down critical computations. In contrast, Perseus can precisely slow down non-critical computations, effectively reducing energy bloat.
6.5. Overhead of Perseus
@@
In the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage. For our A100 workloads, the initial profiling phase increased end-to-end training time by 13 minutes on average, which is negligible overhead for large model training.
@@
The average runtime of the optimization algorithm (§4) across our A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes). For our largest scale emulation (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds, which is shorter because it is sufficient to optimize just one data parallel copy (§4.4). While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time also increases with the scale of the training job. Looking up the optimal energy schedule given $T^{\prime}$ is instant.
7. Related Works
@@
Many recent works focus on enabling and accelerating large model training using 3D parallelism. GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) were the first to introduce pipeline parallelism. Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ) is a large model training framework that provides efficient 3D parallel execution plans tailored for Transformer models. 3D parallelism is considered to be go-to solutions for modern large model development due to strong open-source implementations (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively simple implementation complexity. Finally, Alpa (alpa-osdi22, ) and GSPMD (spmd_socc22, ) provide automatic parallelization for general DNNs. However, energy consumption is not an optimization metric for any of the major large model training frameworks.
@@
A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ) and proposed optimization methods (odpp, ; gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; depo, ; envpipe-atc23, ). Especially, Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on single-GPU training. EnvPipe (envpipe-atc23, ), on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown. However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings. Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm. Perseus is also the first to introduce the notion of energy bloat.
8. Conclusion
We presented Perseus, a software-based energy optimization system for large model training. Perseus builds on the observation that there are computation imbalances at different levels in large model training that causes intrinsic and extrinsic energy bloat, and introduces a principled graph cut-based algorithm that simultaneously reduces both. As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby reducing average power draw as well.
Appendix A Visualizations for Intrinsic Energy Bloat
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 10. Visualization of Perseus’s $T_{\min}$ solution for four stage pipeline workloads on NVIDIA A100 PCIe GPUs. For each workload, the left is running every computation at maximum frequency, and the right is Perseus’s energy schedule that reduces only intrinsic energy bloat without inflating iteration time. Note that these are not real workloads we run in Section 6; real workloads have far more microbatches.
Appendix B Workload Details
B.1. Minimum Imbalance Pipeline Partitioning
@@
(b) NVIDIA A40 GPUs.
Table 7. Imbalance ratio between the longest and the shortest stages for various models. 1.00 would mean perfect balance. Partitions for $N$ stages is a list of $N+1$ numbers, where the numbers represent layer indices. For instance, [0, 6, 12, 19, 25] for GPT-3 1.3B means there are 6, 6, 7, and 5 Transformer layers in each stage, and the final stage also has the language model head.
@@
Arguably, these are one of the most homogeneous large models because they are a stack of identical Transformer [65] encoder or decoder layers. However, the very last layer is the language modeling head, which maps the final features to probabilities over the entire vocabulary. The vocabulary size of GPT-3 is 50k, Bloom 251k, and BERT 31k, which results in a very large linear layer for the last stage. This leads to varying amounts of imbalance and different minimum imbalance partitioning results for each model.
@@
B.2. Alternative Planning Methods
3D parallelism is the go-to solution for large model training, and our minimum imbalance pipeline partitioning method is optimal for 3D parallelism because we implemented a brute force search. More advanced planning methods such as Alpa [77] exist, but for repetitive language model architectures like GPT-3, Alpa equally allocates Transformer layers to each stage [77], resulting in the same stage imbalance caused by the language modeling head. Furthermore, it is hard for tensor parallelism to divide operations infinitely; practically at most degree 8 (within one node) due to collective communication overhead. Thus, in the vast majority of models, there will always be some degree of imbalance between stages due to the minimum granularity of computation time.
B.3. Experiment Parameters
@@
Table 8. Experiment Parameters for 3D parallelism experiments on A40 GPUs. Microbatch size is per-pipeline, and there are two data parallel copies of the same pipeline. Thus, global batch size should be calculated as the product of microbatch size and the number of microbatches times two.
@@
Table 9. Experiment Parameters for eight-stage pipeline parallelism experiments on A40 GPUs. Model variant names are as described in Torch Vision [49] or Huggingface Hub [68].
@@
Table 10. Experiment Parameters for Pipeline Parallelism Experiments on A100 PCIe GPUs. Model variant names are as described in Torch Vision [49] or Huggingface Hub [68].
Tables 89, and 10 list model variant names and experiment parameters for our experiments. Model names and configurations for GPT-3 were taken as is from the original model publication [9]. Especially, model names and configurations for BERT and T5 were directly taken from the Huggingface Hub pretrained model zoo, except for the huge variant of BERT, which we created to have hidden dimension 2048. Wide-ResNet was based on Torch Vision [49] but scaled up following the model’s original publication [74] using its width factor parameter. The unit time parameter $\tau$ was set to 1 ms for all experiments.
Appendix C Pipeline Energy Minimization is NP-hard
The Pipeline Energy Minimization problem is stated again for convenience:
(6) $\min_{F}$ $\textrm{Energy}(F)$
@@ Appendix C Pipeline Energy Minimization is NP-hard
C.1. One Stage Two Frequencies Simplification
@@
C.3. NP-hardness Proof
Theorem C.1.
@@ C.3. NP-hardness Proof
Appendix D Full Details of GetNextPareto
@@ D.1. Generating Capacity DAG
(7) $\min_{I^{+},~{}I^{-}}\quad\sum_{i\in I^{+}}{e_{i}^{+}}-\sum_{i\in I^{-}}{e_{i}^{-}},$
where $e_{i}(t_{i})$ is the exponential function fit to Pareto-optimal (time, energy) measurements for computation $i$, $e_{i}^{+}=e_{i}(t_{i}-\tau)-e_{i}(t_{i})$ is the extra amount of energy needed to speed up $i$ by $\tau$ and $e_{i}^{-}=e_{i}(t_{i})-e_{i}(t_{i}+\tau)$ is the energy saved from slowing down $i$ by $\tau$.
@@
(8) $(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$
@@
D.2. Max Flow Algorithm on the Capacity DAG
@@
5
6
7
8Construct a new graph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding new source and sink nodes $s^{\prime}$ and $t^{\prime}$, edges from $s^{\prime}$ to each node in $V$, edges from each node in $V$ to $t^{\prime}$, and an edge from $t$ to $s$
9
$\triangleright$ Define the capacity $c^{\prime}(e)$ of each edge $e\in E^{\prime}$
10: for $v\in V$ do
11       $c^{\prime}(s^{\prime}v)\leftarrow\sum_{u\in V}l(uv)$
12       $c^{\prime}(vt^{\prime})\leftarrow\sum_{w\in V}l(vw)$
13
@@ 16       $c^{\prime}(uv)\leftarrow u(uv)-l(uv)$
[…]
20
$\triangleright$ Find the max flow on G’
21 $f^{\prime}\leftarrow$ EdmondsKarp ($G^{\prime}$, $c^{\prime}$)
22
$\triangleright$ $G$ has a feasible flow if and only if $G^{\prime}$ has a saturating flow
@@ Appendix E Proof for Polynomial Runtime
$t_{\max}$ and $t_{\min}$ can be constructed by multiplying the number of computations with the average execution time of a computation. Computations are executed with frequencies $f_{\min}$ and $f_{\max}$, respectively, and thus the multipliers of $N$ and $M$ do not cancel out when $t_{\max}-t_{\min}$ is evaluated. Therefore, $t_{\max}-t_{\min}$, and hence $(t_{\max}-t_{\min})/\tau$, is $O(N+M)$. ∎
{listing}{minted}
[]python # In the framework’s pipeline execution engine: from perseus.client import profiler, controller
def train_step(model, dataloader): …
for instuction in pipeline_schedule: if isinstance(instruction, Forward): controller.set_speed(”forward”) profiler.begin(”forward”) # Run forward on microbatch profiler.end(”forward”) elif isinstance(instruction, Backward): controller.set_speed(”backward”) profiler.begin(”backward”) # Run backward on microbatch profiler.end(”backward”) … Perseus client API integration example.
Appendix F Perseus Client Integration Example
Listing E shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
A typical structure of a pipeline execution engine is to have instructions for each distinct operations in the pipeline, including not only forward and backward executions, but also P2P and collective communications, and implement a handler for each instruction. Therefore, framework developers can wrap such handlers with the Perseus client APIs to mark their beginning and end.
Appendix G Iteration Time–Energy Frontiers